OpenStudy (anonymous):

(n-2)log7.6 = log41.7 I know how to input logarithms into the calculator and all that stuff, but I don't know the order of operations when it comes to the (n-2) at the front. How would I add the (n-2) into the problem?

zepdrix (zepdrix):

\[\Large\bf\sf (n-2)\log7.6\quad=\quad \log41.7\]So you want to solve for n, and you just want to make sure you apply your math rules in the correct order?

OpenStudy (anonymous):

Yes.

zepdrix (zepdrix):

We'll start be getting (n-2) alone on the left side. See how we have something multiplying the brackets? So we'll do the inverse in order to isolate the (n-2). We'll start by dividing both sides by log7.6

zepdrix (zepdrix):

\[\Large\bf\sf (n-2)\quad=\quad \frac{\log41.7}{\log7.6}\]Does that step make sense? The log7.6/log7.6 on the left side cancelled out.

OpenStudy (anonymous):

Oh, okay! And then add two, and get about 3.84. Thank you!

zepdrix (zepdrix):

Got it? :) yay!